What are the values of the critical constants Vc , Tc and Pc ? Hence, show that the value of the critical coefficient is
8/3.
1
Expert's answer
2019-10-22T10:22:00-0400

The values of the critical constants:


"T_c=\frac{8a}{27Rb}"

"P_c=\frac{a}{27b^2}"

"V_c=3b"

The value of the critical coefficient is  


"\frac{RT_c}{P_cV_c}=\frac{R(\frac{8a}{27Rb})}{(\frac{a}{27b^2})3b}=\frac{8}{3}"